# Faster Convergence on Heterogeneous Federated Edge Learning: An Adaptive Clustered Data Sharing Approach

### v2

@@ Abstract
Federated Edge Learning (FEEL) emerges as a pioneering distributed machine learning paradigm for the 6G Hyper-Connectivity, harnessing data from the Internet of Things (IoT) devices while upholding data privacy. However, current FEEL algorithms struggle with non-independent and non-identically distributed (non-IID) data, leading to elevated communication costs and compromised model accuracy. To address these statistical imbalances within FEEL, we introduce a clustered data sharing framework, mitigating data heterogeneity by selectively sharing partial data from cluster heads to trusted associates through sidelink-aided multicasting. The collective communication pattern is integral to FEEL training, where both cluster formation and the efficiency of communication and computation impact training latency and accuracy simultaneously. To tackle the strictly coupled data sharing and resource optimization, we decompose the overall optimization problem into the clients clustering and effective data sharing subproblems. Specifically, a distribution-based adaptive clustering algorithm (DACA) is devised basing on three deductive cluster forming conditions, which ensures the maximum sharing yield. Meanwhile, we design a stochastic optimization based joint computed frequency and shared data volume optimization (JFVO) algorithm, determining the optimal resource allocation with an uncertain objective function. The experiments show that the proposed framework facilitates FEEL on non-IID datasets with faster convergence rate and higher model accuracy in a limited communication environment.
@@ I Introduction
In the FL framework, a number of federated rounds are required between clients and servers to collaboratively achieve desired accuracy [4, 5, 6]. Given the nature of deep learning models containing millions of parameters, e.g., the Large Language Model (LLM) GPT-3.5 has 175 billion parameters [7], transmitting such high-dimensional models results in a substantial communication costs. By balancing communication and computation resources of edge devices, Federated Edge learning (FEEL) has been proposed to reduce latency and energy consumption in wireless networks [8]. Yet the training costs are still serious especially when training with non-independent and non-identically distributed (non-IID) data across devices in FEEL. In practice, due to the decentralized and geographically dispersed nature of edge devices, data within a federated learning environment displays significant variations and samples imbalance, i.e., statistical heterogeneity. Extensive research studies [9, 10, 11, 12] have shown that as the degree of heterogeneity increases, the convergence of the FEEL algorithm slows down, and more communication rounds are required for the desired accuracy. Therefore, the time consumption and weak accuracy caused by the non-IID data become the bottleneck of FEEL applied in real-world IoT applications.
@@
To mitigate the communication cost, some other studies try to design resource-efficient FEEL system. These approaches aim to minimize FL loss while optimizing resource efficiency. Unfortunately, these two goals largely conflict with each other, as obtaining a higher-quality model often necessitates increased resource consumption, and vice versa. Chen et al. [25] introduce a joint learning and communication framework that derives a closed-form expression for convergence based on packet error rates, and formulates optimal strategies for user selection and resource allocation to minimize the FEEL loss function. Zhao et al. [11] jointly minimizes the accuracy loss and the cost of data sharing to mitigate the impact of non-IID data, and proposes a computation-efficient algorithm to approach a stationary optimal solution. [26] proposes a resource-efficient method for training a FL model with non-IID data, effectively minimizing cost through an auction approach and mitigating quality degradation through data sharing. [9] analyzes the implicit connection between data distribution and trained models from empirical and mathematical perspectives, and proposes a reinforcement learning-based mechanism for selecting the optimal subset of devices. Therefore, it is a majority challenge for FEEL to accelerate training process in communication and computation meanwhile ensuring model performance with non-IID data.
@@
Although these approaches have attempted to address the statistical heterogeneity issue at model-driven, data-driven, and system design levels, deploying non-IID FL in wireless communication systems still faces challenges. Firstly, model-driven methods demand additional computing resources and show marginal improvements. The increased complexity of local training adds to the difficulty of applying FL in resource-constrained wireless networks. Much work [26, 5, 11, 23] demonstrate that data sharing is a straightforward and efficient method for directly mitigating model quality degradation. Secondly, most data sharing methods work under the strongly assumption of a publicly available proxy data source on the server side [21, 26]. However, building a pre-processed and labeled public dataset is a time-consuming and expensive task. Evaluating the quality of the shared dataset is also critical, since it directly ensures performance gains. Lastly, the costs and privacy concerns of data sharing cannot be ignored. Exchanging data unavoidably raises information leakage, additional latency, and energy consumption. Specifically, when an excessive amount of data is transferred, the costs and privacy risks may outweigh the benefits of sharing data.
@@
Based on the above analysis, this work proposes a data sharing framework that allows partial data exchanged among trusted user groups to reduce the degree of data heterogeneity. In the real-world scenario, there are certain devices with high-quality data that can accelerate FEEL training. By the aid of the high-bandwidth mmWave multicasting technology [27], these nodes can share the private data with trustable neighbors in close proximity communication (i.e., sidelink) directly without going through the base station (BS). These sidelink transmissions boost network ulity while saving spectrum resource since it enables the simultaneous delivery of content to multiple users through a single transmission [28]. By selectively exchanging data with trustworthy and dependable partners via sidelink-aided technology, communication overhead and privacy concerns can be effectively regulated. Moreover, the goals of this paper are not only to mitigate the statistical heterogeneity issue, but also to accelerate FEEL training by exploiting data and computility of edge network. The key contributions of this paper are summarized as follows:
@@
  • Clustered data sharing FEEL framework for eliminating statistical heterogeneity: Inspired by the valid results of data-driven methods on non-IID FL, we propose a clustered data sharing framework which can evaluate and eliminate the impacts of statistical heterogeneity. Sidelink-aided multicasting is used as an enabling technology of wireless network for FL to achieve faster training and higher accuracy. Moreover, this framework can serve as a data preparation procedure applicable to model-driven non-IID FL algorithms.
@@
  • Fast convergence with joint communication, computation configuration: An optimization problem is formulated to minimize the overall delay of the FL task with clustered data sharing by jointly considering communication, computation, and privacy. Here, devices are allowed to trustfully cluster and exchange data to accelerate FEEL training convergence. The shared data volume and computed frequency are jointly optimized to achieve trade-off among sharing delay, communication delay, and computation delay over the training rounds.
@@
  • Performance evaluation from various perspectives: Our proposed approaches are evaluated by applying classic image classification tasks to the MNIST and CIFAR-10 datasets. The clustered data sharing FL with DACA can effectively improve the accuracy in different non-IID cases and significantly reduce the total training task delay. Moreover, The experimental results shows that the proposed JFVO algorithm can convergence in different communication states.
The rest of this paper is structured as follows: Section II explores and analyzes the FL performance with heterogeneous data. Section III presents a clustered data sharing FL framework over wireless networks. Section IV formulates a problem aiming to minimize delays while mitigating the impact of non-IID data, and a joint optimization algorithm is designed to make a good tradeoff between the performance improvement and the cost. Section V provides the numerical results, and finally conclusions are drawn in Section VI.
@@ II-C Preliminary Experiments and Analysis
Figure 1: Train loss on clustered data sharing for non-IID FL. (a) For IID setting, each user is randomly assigned a uniform distribution over all classes. For non-IID setting, the proportion of labels of among users are different, leading to varying ${\bar{D}_{{\text{EMD}}}}$. (b) The shared data volume is quantified as $\alpha$, representing the proportion of sharing data volume in its local data volume. (c) The users are randomly divided into clusters, where each cluster consists of cluster heads to transmit data and cluster members to receive data.
To mitigate the divergence between the local and global distributions, one possible approach is to share data among the participating clients. Traditional data sharing schemes involve a centralized server sharing a portion of the global data, which is made available to all clients. However, this approach requires the creation of a publicly accessible proxy data source, which can be impractical or pose privacy concerns. Instead of centralized sharing, we randomly select clients to exchange local data within clusters. The preliminary experiments between this scheme and the standard FedAvg algorithm reveal that the convergence of the training process is impacted by three main factors.
Degree of heterogeneity: reflects the skewness value of all local data distributions in the overall system, i.e., ${\bar{D}_{{\text{EMD}}}}$. To further investigate the relationship between degree of heterogeneity and FL training performance, we lay out the experiment on the training loss of non-IID FL with different ${\bar{D}_{{\text{EMD}}}}$ settings. As shown in Fig. 1a, it is clear from the result that with the decrease in data heterogeneity, the convergence rate of the FL training is significantly improved, and the loss value of training becomes lower.
@@ III System Model
Figure 2: Clustered data sharing framework for FEEL.
III-A Clustered Data Sharing FEEL Framework
@@
The data preparing stage: involves a data augmentation procedure before the FEEL training. In this stage, all users are clustered with respect to data distribution, communication states, and privacy requirements, where each cluster is defined by a cluster $m$ and its corresponding cluster members ${{\cal C}_{m}}$. Then the cluster heads, denoted as the set ${\cal M}$, share the subsets of their data with cluster members through reliable multicast communication. Following clustered data exchange, each node can compute its local parameters with the mixed samples in the next stage. It is noted that each node cannot participate in more than one cluster to prevent conflicts, i.e., ${{\cal C}_{m}}\cap{{\cal C}_{l}}=\emptyset,\forall m,l\in{\mathcal{M}},m\neq l$. With a copy of shared data, the local data volume on device $k$ is then determined by the following expression:
@@
The FEEL training stage: is to train a global model via collaborations between the edge server and multiple clients, where the global model is obtained by minimizing the following loss function:
@@
After model aggregation, $w_{g}^{t}$ is fed back to the devices, and the FEEL above training procedure repeats for multiple rounds until convergence.
@@ III-B Clustered Data Sharing Model
${\gamma\left({{d_{m,c}}}\right)=\sum\limits_{i=0}^{3}{\frac{{{C_{i}}A_{i}^{-1}d_{m,c}^{-{\zeta_{i}}}{\kappa_{i}}\left({{d_{m,c}}}\right){S_{F,i}}{U_{F,i}}}}{{{N_{0}}W+I}}}},$ (8)
where ${C_{i}}={P_{m}}{G_{m}}{G_{c}}$, and $P_{m}$ is the transmit power, $G_{m}$ and $G_{c}$ are the antenna array gains at cluster head $m$ and member $c$. $d_{m,c}$ denotes the three-dimensional (3D) distance between cluster head $m$ and member $c$. ${\kappa_{i}},i\in\left\{{0,1,2,3}\right\}$ is the state probabilities of four states: non-line-of-sight (nLoS) with blocked, line-of-sight (LoS) with blocked, nLOS with non-blocked, LOS with non-blocked. $A_{i}$ and $\zeta_{i}$ are propagation coefficients of the path loss, which corresponding to the introduced states. ${S_{F,i}}$ and ${U_{F,i}}$ are state-dependent shadow and small-scale fading, $I$ is the interference. In the urban microcell (UMi) scenario, we set these coefficient values according to the 3GPP standards [36]. With multicasting, the transmission rate of data sharing within one cluster is determined by the worst sidelink, i.e.,
@@
where $a$ denotes bits of each data sample. The energy consumption of one-shot data sharing can be ignored compared to the multi-round FEEL training process.
III-C FEEL Transmission and Computation Model
As introduced in Section III-A, the FEEL training stage comprises four steps in each round, with each step incurring either communication or computational cost.
@@ III-C FEEL Transmission and Computation Model
According to the system model, the FEEL latency is composed of three parts: downloading, computation, and uploading delay, and all clients have to complete the synchronization of the broadcast before proceeding with any local update. Hence, the total delay of one FEEL round with the users and the BS jointly participating can be expressed as
@@
Due to the fact that communication rounds are independent, it suffices to consider the energy consumption for an arbitrary round without loss of generality [37]. Therefore, the total energy consumption for a single FEEL round, considering both computation and uploading energy, can be expressed as
@@ IV-A Problem Formulation and Decomposition
To maximize the effectiveness of data sharing in accelerating FEEL training, we formulate an optimization framework to minimize the total delay for training the whole FEEL algorithm. This minimization problem joints the design of the clustering strategy and resource optimization (i.e., shared data volume and computed frequency) for each cluster while guaranteeing privacy as follows:
@@
where ${{\cal C}}{\rm{=}}\left\{{\left.{{{{\cal C}}_{1}}\ldots,{{{\cal C}}_{m}},\ldots}\right\}}\right.$ and ${{{\cal C}}_{m}}$ is the set of cluster members connected with the cluster head $m$. $\boldsymbol{F}=\left[{{f_{1}},\ldots,{f_{K}}}\right]$ is a computed frequency vector with $K$ being the total number of devices. ${\boldsymbol{N}}^{s}=\left[{n_{1}^{s},\ldots,n_{m}^{s},\ldots}\right]$ is a vector of sharing data volume from the cluster heads. $T$ is the number of global round needed to attain the target accuracy $\theta_{{\text{th}}}$. Constraint (20a) is the preset accuracy requirement for FL training. (20b) is the constraint for exclusive cluster forming. Constraint (20c) restricts the privacy requirement for data sharing between cluster heads and cluster members. The shared data volume should satisfy constraint (20d). $f_{\max}$ is the maximum local computation capacity, and constraint (20e) presents the frequency limits of all devices. Constraint (20f) is the energy consumption requirement of FEEL training at each round. Accordingly, the multi-constraints involved objective achieve the joint optimization of communication, computation and learning performance by exploiting tradeoff of the data and computility resource at the edge.
@@
In order to solve problem $\mathcal{P}0$, we adopt a decomposition approach by recognizing that the FEEL training process can only commence once the clustering process has been completed. As such, we break down the primal problem into the following two subproblems, based on the sequential order of the clustering stage and the FEEL training stage.
Subproblem 1: Efficient Cluster Forming Strategy with Precision and Privacy Constraints. The optimal clustering strategy is designed to encompass the entire clustering and FEEL training stages, with the objective of minimizing the overall delay, i.e.,

### v3

@@ Abstract
Federated Edge Learning (FEL) emerges as a pioneering distributed machine learning paradigm for the 6G Hyper-Connectivity, harnessing data from the IoT devices while upholding data privacy. However, current FEL algorithms struggle with non-independent and non-identically distributed (non-IID) data, leading to elevated communication costs and compromised model accuracy. To address these statistical imbalances, we introduce a clustered data sharing framework, mitigating data heterogeneity by selectively sharing partial data from cluster heads to trusted associates through sidelink-aided multicasting. The collective communication pattern is integral to FEL training, where both cluster formation and the efficiency of communication and computation impact training latency and accuracy simultaneously. To tackle the strictly coupled data sharing and resource optimization, we decompose the optimization problem into the clients clustering and effective data sharing subproblems. Specifically, a distribution-based adaptive clustering algorithm (DACA) is devised basing on three deductive cluster forming conditions, which ensures the maximum sharing yield. Meanwhile, we design a stochastic optimization based joint computed frequency and shared data volume optimization (JFVO) algorithm, determining the optimal resource allocation with an uncertain objective function. The experiments show that the proposed framework facilitates FEL on non-IID datasets with faster convergence rate and higher model accuracy in a resource-limited environment.
@@ I Introduction
In the FL framework, a number of federated rounds are required between clients and servers to collaboratively achieve desired accuracy [4, 5, 6]. Given the nature of deep learning models containing millions of parameters, e.g., the Large Language Model (LLM) GPT-3.5 has 175 billion parameters [7], transmitting such high-dimensional models results in a substantial communication costs. By balancing communication and computation resources of edge devices, Federated Edge learning (FEL) has been proposed to reduce latency and energy consumption in wireless networks [8]. Yet the training costs are still serious especially when training with non-independent and non-identically distributed (non-IID) data across devices in FEL. In practice, due to the decentralized and geographically dispersed nature of edge devices, data within a federated learning environment displays significant variations and samples imbalance, i.e., statistical heterogeneity. Extensive research studies [9, 10, 11, 12] have shown that as the degree of heterogeneity increases, the convergence of the FEL algorithm slows down, and more communication rounds are required for the desired accuracy. Therefore, the time consumption and weak accuracy caused by the non-IID data become the bottleneck of FEL applied in real-world IoT applications.
@@
To mitigate the communication cost, some other studies try to design resource-efficient FEL system which minimize FL loss while optimizing resource efficiency. Chen et al.[25] introduce a joint learning and communication framework considering resource allocation and user selection. [26] proposes a multi-agent reinforcement learning-based framework for optimizing device selection and resource allocation in distributed industrial IoT networks. [27] proposes a Cluster-based Parallel Split Learning (CPSL) scheme to reduce training latency through a combination of parallel intra-cluster and sequential inter-cluster learning. These approaches These methods optimize communication resources in FEL but not considering the impact of distrusted data distribution on training performance. [9] analyzes the implicit connection between data distribution and FEL training, and proposes a reinforcement learning-based mechanism for user selection. Zhao et al. [11] jointly minimizes the accuracy loss, delay and energy of FEL with data sharing to mitigate the impact of non-IID data. [28] proposes an auction approach to minimize the training and sharing cost in non-IID FEL. These frameworks rely on uploading local data to the server, which raises concerns about privacy leakage.
@@
Although model-driven, data-driven, and system-level design approaches have attempted to address the statistical heterogeneity issue, there remain significant technical limitations deserving further research. Firstly, model-driven methods demand additional computing resources and show unclear improvements relying on specific hyperparameters. The increased complexity of local training adds to the difficulty of applying FL in resource-constrained wireless networks. Secondly, much work [28, 5, 11, 23] demonstrate that data sharing is a straightforward and efficient method for directly mitigating model quality degradation. However, most methods work under the strongly assumption of a publicly available proxy data source on the server side [21, 28]. Building a pre-processed and labeled public dataset is a time-consuming and expensive task. Lastly, the communication costs and privacy concerns of data sharing cannot be ignored. Exchanging data unavoidably raises information leakage, additional latency, and energy consumption. To address these limitations, we adopt direct data exchange among trusted users using high-speed mmWave multicasting technology. This removes the dependency on pre-prepared datasets or hyperparameters, reducing both privacy risks and communication overhead.
@@
Based on the above analysis, this work proposes a data sharing framework that allows partial data exchanged among trusted user groups to reduce the degree of data heterogeneity. In the real-world scenario, there are certain devices with high-quality data that can accelerate FEL training. By the aid of the high-bandwidth mmWave multicasting technology [29], these nodes can share the private data with trustable neighbors in close proximity communication (i.e., sidelink) directly without going through the base station (BS). These sidelink transmissions boost network ulity while saving spectrum resource since it enables the simultaneous delivery of content to multiple users through a single transmission [30]. By selectively exchanging data with trustworthy and dependable partners via sidelink-aided technology, communication overhead and privacy concerns can be effectively regulated. Moreover, the goals of this paper are not only to mitigate the statistical heterogeneity issue, but also to accelerate FEL training by exploiting data and computility of edge network. The key contributions of this paper are summarized as follows:
@@
  • Clustered data sharing FEL framework for eliminating statistical heterogeneity: We propose a clustered data sharing framework which can evaluate and eliminate the impacts of statistical heterogeneity in FEL. Sidelink-aided multicasting is used as an enabling technology of wireless network for FL to achieve faster training and higher accuracy. By appropriately modifying the data distribution function, the framework can be generalized to other non-IID causes, demonstrating its robustness and scalability. Moreover, this framework can also serve as a data preparation procedure applicable to model-driven non-IID FL algorithms.
@@
  • Fast convergence with joint communication, computation configuration: An optimization problem is formulated to minimize the overall delay of the FL task with clustered data sharing by jointly considering communication, computation, and privacy. Here, devices are allowed to trustfully cluster and exchange data to accelerate FEL training convergence. The shared data volume and computed frequency are jointly optimized to achieve trade-off among sharing delay, communication delay, and computation delay over the training rounds.
@@
  • Performance evaluation from various perspectives: Our proposed approaches are evaluated by applying classic image classification tasks to the MNIST and CIFAR-10 datasets. The clustered data sharing FL with DACA can effectively improve the accuracy in different non-IID cases and significantly reduce the total training task delay. Moreover, The experimental results shows that the proposed JFVO algorithm can converge in different communication states.
@@ II-C Preliminary Experiments and Analysis
To mitigate the divergence between the local and global distributions, one possible approach is to share data among the participating clients. Traditional data sharing schemes involve a centralized server sharing a portion of the global data, which is made available to all clients. However, this approach requires the creation of a publicly accessible proxy data source, which can be impractical or pose privacy concerns. Instead of centralized sharing, we randomly select clients to exchange local data within clusters. The skewness of the overall system is quantified using the average EMD ${\bar{D}_{\text{EMD}}}$. EMD is widely recognized as an effective metric for quantifying the statistical heterogeneity of data distributions in FL [9]. To ensure fairness, the average EMD used in our work is weighted by the data volume of each user. Specifically, if users with significant deviations in data distribution have smaller dataset sizes, their influence on the average EMD is minimal. Conversely, users with larger dataset sizes and substantial distribution deviations exert a greater influence, resulting in a higher average EMD value. The preliminary experiments between this scheme and the standard FedAvg algorithm reveal that the convergence of the training process is impacted by three main factors.
@@
Figure 1: Train loss on clustered data sharing for non-IID FL. (a) For IID setting, each user is randomly assigned a uniform distribution over all classes. For non-IID setting, the proportion of labels of among users are different, leading to varying ${\bar{D}_{{\text{EMD}}}}$ [^2] Please note that we assume all users have equal data volume, with only the proportion of labels changes, result in different average EMD.. (b) The shared data volume is quantified as $\alpha$, representing the proportion of sharing data volume in its local data volume. (c) The users are randomly divided into clusters, where each cluster consists of cluster heads to transmit data and cluster members to receive data.
Degree of heterogeneity: To further explore the relationship between data heterogeneity and FL training performance, we conduct experiments on training accuracy under varying $\bar{D}_{\text{EMD}}$ settings in label distribution skew case. As shown in Fig. 1a, the results clearly indicate that reducing data heterogeneity significantly enhances the convergence rate of FL training, leading to higher model accuracy. These findings demonstrate that a higher $\bar{D}_{\text{EMD}}$ slows the accuracy improvement per communication round, requiring more rounds to reach the same accuracy level as configurations with lower $\bar{D}_{\text{EMD}}$.
@@ III System Model
Figure 2: Clustered data sharing framework for FEL.
III-A Clustered Data Sharing FEL Framework
@@
The data preparing stage: involves a data augmentation procedure before the FEL training. In this stage, all users are clustered with respect to data distribution, communication states, and privacy requirements, where each cluster is defined by a cluster $m$ and its corresponding cluster members ${{\cal C}_{m}}$. Then the cluster heads, denoted as the set ${\cal M}$, share the subsets of their data with cluster members through reliable multicast communication. Following clustered data exchange, each node can compute its local parameters with the mixed samples in the next stage. It is noted that each node cannot participate in more than one cluster to prevent conflicts, i.e., ${{\cal C}_{m}}\cap{{\cal C}_{l}}=\emptyset,\forall m,l\in{\mathcal{M}},m\neq l$. With a copy of shared data, the local data volume on device $k$ is then determined by the following expression:
@@
The FEL training stage: is to train a global model via collaborations between the edge server and multiple clients, where the global model is obtained by minimizing the following loss function:
@@
After model aggregation, $w_{g}^{t}$ is fed back to the devices, and the FEL above training procedure repeats for multiple rounds until convergence.
@@ III-B Clustered Data Sharing Model
${\gamma\left({{d_{m,c}}}\right)=\sum\limits_{i=0}^{3}{\frac{{{C_{i}}A_{i}^{-1}d_{m,c}^{-{\zeta_{i}}}{\kappa_{i}}\left({{d_{m,c}}}\right){S_{F,i}}{U_{F,i}}}}{{{N_{0}}{B_{m}^{s}}+I}}}},$ (8)
where ${C_{i}}={P_{m}}{G_{m}}{G_{c}}$, and $P_{m}$ is the transmit power, $G_{m}$ and $G_{c}$ are the antenna array gains at cluster head $m$ and member $c$. $d_{m,c}$ denotes the three-dimensional (3D) distance between cluster head $m$ and member $c$. ${\kappa_{i}},i\in\left\{{0,1,2,3}\right\}$ is the state probabilities of four states: non-line-of-sight (nLoS) with blocked, line-of-sight (LoS) with blocked, nLOS with non-blocked, LOS with non-blocked. $A_{i}$ and $\zeta_{i}$ are propagation coefficients of the path loss, which corresponding to the introduced states. ${S_{F,i}}$ and ${U_{F,i}}$ are state-dependent shadow and small-scale fading. $N_{0}$ is the noise power spectral density, $B_{m}^{s}$ is the bandwidth of multicasting in cluster $m$ and $I$ is the interference. In the urban microcell (UMi) scenario, we set these coefficient values according to the 3GPP standards [38]. With multicasting, the transmission rate of data sharing within one cluster is determined by the worst sidelink, i.e.,
@@
where $a$ denotes bits of each data sample. The energy consumption of one-shot data sharing can be ignored compared to the multi-round FEL training process.
III-C FEL Transmission and Computation Model
As introduced in Section III-A, the FEL training stage comprises four steps in each round, with each step incurring either communication or computational cost.
@@ III-C FEL Transmission and Computation Model
According to the system model, the FEL latency is composed of three parts: downloading, computation, and uploading delay, and all clients have to complete the synchronization of the broadcast before proceeding with any local update. Hence, the total delay of one FEL round with the users and the BS jointly participating can be expressed as
@@
Due to the fact that communication rounds are independent, it suffices to consider the energy consumption for an arbitrary round without loss of generality [39]. Therefore, the total energy consumption for a single FEL round, considering both computation and uploading energy, can be expressed as
@@ IV-A Problem Formulation and Decomposition
To maximize the effectiveness of data sharing in accelerating FEL training, we formulate an optimization framework to minimize the total delay for training the whole FEL algorithm. This minimization problem joints the design of the clustering strategy and resource optimization (i.e., shared data volume and computed frequency) for each cluster while guaranteeing privacy as follows:
@@
where ${{\cal C}}{\rm{=}}\left\{{\left.{{{{\cal C}}_{1}}\ldots,{{{\cal C}}_{m}},\ldots}\right\}}\right.$ and ${{{\cal C}}_{m}}$ is the set of cluster members connected with the cluster head $m$. $\boldsymbol{F}=\left[{{f_{1}},\ldots,{f_{K}}}\right]$ is a computed frequency vector with $K$ being the total number of devices. ${\boldsymbol{N}}^{s}=\left[{n_{1}^{s},\ldots,n_{m}^{s},\ldots}\right]$ is a vector of sharing data volume from the cluster heads. $T$ is the number of global round needed to attain the target accuracy $\theta_{{\text{th}}}$. Constraint (20a) is the preset accuracy requirement for FL training. (20b) is the constraint for exclusive cluster forming. Constraint (20c) restricts the privacy requirement for data sharing between cluster heads and cluster members. The shared data volume should satisfy constraint (20d). $f_{\max}$ is the maximum local computation capacity, and constraint (20e) presents the frequency limits of all devices. Constraint (20f) is the energy consumption requirement of FEL training at each round. Accordingly, the multi-constraints involved objective achieve the joint optimization of communication, computation and learning performance by exploiting tradeoff of the data and computility resource at the edge.
@@
In order to solve problem $\mathcal{P}0$, we adopt a decomposition approach by recognizing that the FEL training process can only commence once the clustering process has been completed. As such, we break down the primal problem into the following two subproblems, based on the sequential order of the clustering stage and the FEL training stage.
Subproblem 1: Efficient Cluster Forming Strategy with Precision and Privacy Constraints. The optimal clustering strategy is designed to encompass the entire clustering and FEL training stages, with the objective of minimizing the overall delay, i.e.,